OpenStudy (anonymous):

The slope of a line is -, and the y-intercept is 10/3. What is the equation of the line written in general form? A) 10x + 3y - 1 = 0 B) x + 3y + 10 = 0 C) x + 3y - 10 = 0

OpenStudy (anonymous):

y=-1/5x +10/3

OpenStudy (anonymous):

is it A?

OpenStudy (tkhunny):

Have you seen the Slope-Intercept Form?

OpenStudy (anonymous):

slope is - what?? suppose to be -1/5

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yeah

OpenStudy (quietus):

Answer is C

OpenStudy (anonymous):

YOu cannot determine the slope with the info given. Put each equation in the form y=mx+b...

OpenStudy (quietus):

Follow y=mx+c, y=x/3+10/3

OpenStudy (quietus):

c = y intercept.

OpenStudy (anonymous):

=> y=-1/5 x +10/3

OpenStudy (quietus):

For A & B, the y intercept is not 10/3.

OpenStudy (tkhunny):

Anytime you are given the slope (or a hint about the slope) and the y-intercept, the Slope-Intercept is normally the way to go. A) 10x + 3y - 1 = 0 B) x + 3y + 10 = 0 C) x + 3y - 10 = 0 Move the constant with subtraction. A) 10x + 3y = 1 B) x + 3y = -10 C) x + 3y = 10 Move the x-term with subtraction. A) 3y = -10 x + 1 B) 3y = -x - 10 C) 3y = -x + 10 Isolate the the 'y' and produce the final Slope-Intercept Form. A) y = -(10/3) x + 1/3 B) y = -(1/3)x - 10/3 C) y = -(1/3)x + 10/3 Identify the parts A) y = -(10/3) x + 1/3 <== Slope: -10/3 and y-intercept 1/3 B) y = -(1/3)x - 10/3 <== Slope: -1/3 and y-intercept -10/3 C) y = -(1/3)x + 10/3 <== Slope: -1/3 and y-intercept 10/3 Find the one that fills all the requirements. Be deliberate and systematic. No need to guess or wonder.